# The Road to Artificial SuperIntelligence: A Comprehensive Survey of Superalignment

### v1

@@ Abstract
The emergence of large language models (LLMs) has sparked the possibility of about Artificial Superintelligence (ASI), a hypothetical AI system surpassing human intelligence. However, existing alignment paradigms struggle to guide such advanced AI systems. Superalignment, the alignment of AI systems with human values and safety requirements at superhuman levels of capability aims to addresses two primary goals—scalability in supervision to provide high-quality guidance signals and robust governance to ensure alignment with human values. In this survey, we examine scalable oversight methods and potential solutions for superalignment. Specifically, we explore the concept of ASI, the challenges it poses, and the limitations of current alignment paradigms in addressing the superalignment problem. Then we review scalable oversight methods for superalignment. Finally, we discuss the key challenges and propose pathways for the safe and continual improvement of ASI systems. By comprehensively reviewing the current literature, our goal is provide a systematical introduction of existing methods, analyze their strengths and limitations, and discuss potential future directions.
@@ 1 Introduction
Figure 1: Challenges in the development of ASI from the perspectives of supervision and governance. While supervision perspective focuses on providing high-quality guidance signals for enhancing system competence, governance perspective emphasizes aligning the behavior of ASI systems with human values to prevent harmful outcomes.
@@
Language models have shown improved capabilities as their scale increases, a phenomenon known as emergent abilities Wei et al. (2022a); Brown et al. (2020); Kaplan et al. (2020); Nam et al. (2024); Hu et al. (2024). This scaling has led to the development of large language models (LLMs), including proprietary models such as GPT-o1 OpenAI (2024), Claude 3.5 Anthropic (2024), and Gemini Ultra Team et al. (2024), as well as open-source models like Llama 3.2 Meta (2024), Mixtral Mistral (2024); Jiang et al. (2024), and Qwen 2.5 Team (2024) variants.
@@
The emergent abilities of LLMs have enabled significant advancements across a wide range of tasks. These include natural language understanding Brown et al. (2020), reasoning Wei et al. (2022b); Rein et al. (2024), code generation Chen et al. (2021); Liu et al. (2024), and multilingual translation Shi et al. (2022). Furthermore, LLMs have demonstrated surprising breakthroughs, such as partially passing the Turing test Jones and Bergen (2024) and achieving high accuracy on hard mathematical problems Rein et al. (2024). Recent research indicates that these advancements in LLMs have sparked a shift from artificial narrow intelligence (ANI) to the possibility of artificial general intelligence (AGI), which has not yet been achieved, making the idea of artificial superintelligence (ASI) a potential future scenario Pohl (2015); Batin et al. (2017). ANI refers to the AI systems currently in use, which are designed to perform specific tasks, such as LLMs for healthcare Huang et al. (2024) and software engineering Fan et al. (2023), but they lack the general cognitive abilities of humans. In contrast, AGI represents a theoretical AI model that show human-level capabilities, including reasoning, learning, and adaptability across various domains Bubeck et al. (2023); Fei et al. (2022); Goertzel (2014); Pohl (2015); Batin et al. (2017). The developement in LLMs, particularly their ability to generalize knowledge and exhibit emergent behaviors, has led to greater optimism that achieving AGI is becoming increasingly realistic Bubeck et al. (2023). This development also draws attention to ASI, a hypothetical future stage of AI where machines would not only surpass human intelligence in all areas but also have advanced cognitive functions and sophisticated thinking abilities Mucci and Stryker (2023); Hughes et al. (2024). Considering the potential to achieve ASI in the future, it is essential to envision and prepare for the ASI era.
@@
However, the development of ASI poses challenges, particularly due to the lack of guidance signals when humans cannot supervise AI systems with labeled data. Figure 1 illustrates the challenges in scaling high-quality guidance signals for ASI. Superalignment is defined as “the process of supervising, controlling, and governing artificial superintelligence systems” Jonker and McGrath (2024); Jan Leike (2023) and encompasses two sub-goals from the perspectives of supervision and governance. Achieving these goals requires an iterative and co-evolving process.
From a supervision perspective, the goal is to build high-quality guidance signals to enhance the model’s usefulness. Traditional training methods, such as reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), face scalability issues once model begin to surpass human intelligence. This creates a critical bottleneck, the inability to provide high-quality guidance signals that improve the system’s helpfulness or competence.
@@
From a governance perspective, the goal is to effectively regulate the model’s behavior to ensure it remains harmless and aligned with human values. Without robust governance, ASI could behave in unpredictable, harmful, or even catastrophic. For instance, poorly aligned ASI systems might pursue goals that inadvertently disempower humanity or create existential risks, including the threat of human extinction. To address this, it is essential not only to constrain an ASI’s behavior but also to ensure that these systems operate in a manner that is inherently safe and aligned with ethical principles.
This process involves developing methods to guide advanced AI systems, ensuring they remain beneficial and safe. Superalignment enables developing value-aligned ASI, maximizing benefits for humanity.
@@
Current alignment paradigms, such as reinforcement learning with human feedback (RLHF) Christiano et al. (2017); Ouyang et al. (2022); Gulcehre et al. (2023), supervised fine-tuning (SFT) Zhou et al. (2023); Rafailov et al. (2023), and in-context learning (ICL) Gou et al. (2024); Xu et al. (2023), have shown promise in aligning LLMs with human values and goals Wang et al. (2024). RLHF leverages human feedback to train a reward model, which generates rewards that help align the language model with human preferences. In contrast, SFT eliminates the reinforcement reward model and directly fine-tunes the language model using human feedback. Meanwhile, ICL aligns the output of LLMs with human values by constraining their responses during inference without the need for additional training. Scalable oversight, a concept derived from superalignment, aims to develop scalable, high-quality supervision signals that can guide AI systems beyond human capabilities while ensuring alignment with human values and goals Amodei et al. (2016); Taylor et al. (2016); Olah (2016). However, despite advancements in RLHF, SFT, and ICL, these paradigms struggle to achieve the goal of superalignment.
@@
From a supervision perspective, RLHF is limited by the scalability and reliability of human feedback. As models grow in complexity, collecting feedback that is accurate, consistent, and comprehensive becomes increasingly difficult, leaving gaps in guidance. Similarly, SFT faces challenges in curating and annotating datasets that are sufficiently diverse and representative to fine-tune ASI effectively. ICL, while bypassing training entirely, relies on carefully crafted prompts during inference, which are inherently constrained in scalability and expressiveness. From a governance perspective, RLHF and SFT depend heavily on the quality and intent of human supervision, raising concerns about biases and inconsistencies in feedback. The potential for misaligned human oversight poses risks, particularly as models exceed human capabilities. ICL, on the other hand, requires governance structures to ensure that prompt constraints align with broader human values, a process that is difficult to standardize and enforce. Additionally, the risk of deception by advanced AI systems further complicates governance Yang et al. (2024a). Thus, the lack of scalable, high-quality, and diverse supervision signals remains a bottleneck in achieving superalignment. To tackle superalignment challenges, tailored methods must be designed. As models advance towards ASI, it is crucial to enhance their capabilities during development while steering their values and goals to maximize benefits for human society.
@@
Our objective is to provide an overview of the superalignment problem, including its origins, historical context, and definition, as well as the current state of superalignment research and potential solutions. In Section 2, we introduce the concept of ASI, the challenges it poses, and the limitations of current alignment paradigms in addressing the superalignment problem. In Section 3, we analyze current scalable oversight techniques and discuss the specific challenges associated with each method. In Section 4, we summarize the key challenges of scalable oversight and explore potential pathways to address these challenges.
@@ 2.1 Terminology
Artificial Narrow Intelligence (ANI) or “weak AI,” refers to AI systems designed to excel in a single specific task $T$. These systems operate equal or below human-level performance in the task and lack generalization capabilities. This can be formalized as $\mathcal{A}_{\text{ANI}}(T)\leq\mathcal{H}(T)$, where $\mathcal{A}_{\text{ANI}}(T)$ represents the performance of the ANI system on task $T$, and $\mathcal{H}(T)$ represents human-level performance for the same task and domain. Examples of ANI systems include language translation systems and game-playing AI such as chess Mucci and Stryker (2023) and Go Silver et al. (2016).
Artificial General Intelligence (AGI) refers to a theoretical leap beyond ANI. AGI systems would possess human-level intelligence and show general-purpose capabilities across a wide range of tasks and domains Bubeck et al. (2023); Fei et al. (2022); Goertzel (2014); Pohl (2015); Batin et al. (2017). AGI systems generalize knowledge and adapt to new challenges. This can be formalized as $\mathcal{A}_{\text{AGI}}(T_{i})=\mathcal{H}(T_{i}),\quad\forall\,T_{i}$, where $\mathcal{A}_{\text{AGI}}(T_{i})$ represents the AGI system’s performance on task $T_{i}$ with a potentially infinite set of tasks $i\to\infty$, and $\mathcal{H}(T_{i})$ represents human-level performance in the same tasks and domains.
Artificial Superintelligence (ASI) refers to AI systems that surpass human intelligence ($\mathcal{H}$) in all tasks and domains with exceptional thinking skills Mucci and Stryker (2023); Hughes et al. (2024). ASI systems demonstrate superior reasoning, creativity, and adaptability, formally described as $\mathcal{A}_{\text{ASI}}(T_{i})\gg\mathcal{H}(T_{i}),\quad\forall\,T_{i}$, where $\gg$ indicates that the ASI system outperforms human intelligence in quality of results.
Although ASI is currently far from realization, addressing the challenges of superalignment is both timely and critical for mitigating catastrophic risks Pueyo (2018). Proactive preparation is essential, as the emergence of ASI could outpace our ability to respond, leaving no room for correction. Furthermore, supervision and governance must co-evolve as interactive processes, enhancing AI capabilities while simultaneously mitigating associated risks.
@@ 2.2 Superalignment
Scalable oversight is formally defined as “the process of ensuring that a given AI system adheres to aspects of its objectives that are too costly or impractical to evaluate frequently during training” Amodei et al. (2016). This concept shows the need for methods that provide reliable, high-quality supervision signals, even as AI systems evolve beyond human intelligence. Superalignment, as a specific case of scalable oversight, focuses on aligning ASI with human values and intentions while also enhancing its capability to effectively achieve its objectives.
Formally, superalignment is defined as “the process of supervising, controlling, and governing artificial superintelligence systems” Jonker and McGrath (2024) with the goal of developing methods that provide ASI with high-quality supervision signals that guide its development and behavior, even in the absence of direct human oversight or traditional feedback mechanisms such as human-labeled data or RLHF Amodei et al. (2016); Christiano et al. (2018). As AI capabilities grow, these traditional methods become increasingly insufficient due to the escalating cost and expertise required for human supervision.
Since the emergence of large-scale models and the increase in their capabilities, superalignment has gained significant attention as a means to ensure that AI systems remain aligned with human values while maintaining its capability. Scalable oversight techniques, such as sandwiching and weak-to-strong generalization (W2SG), have emerged as promising approaches to address the challenges of governing ASI.
@@ 2.3.2 Methods
Scalable oversight has emerged as a promising solution to the challenge of supervising ASI. It enables the efficient alignment of AI systems with complex objectives that would otherwise be too costly or difficult to evaluate frequently. Various scalable oversight techniques have been proposed, including iterated distillation and amplification (IDA) Christiano et al. (2018); Charbel-Raphaël (2023), recursive reward modeling (RRM) Leike et al. (2018), and cooperative inverse reinforcement learning Hadfield-Menell et al. (2016). These approaches serve as the basis for current scalable oversight methods, such as weak-to-strong generalization, debate, reinforcement learning from AI feedback, and sandwiching Burns et al. (2023); Irving et al. (2018); Bai et al. (2022); Bowman et al. (2022). Figure 2 illustrates the current scalable oversight techniques and their methodologies.
@@
  • Reinforcement Learning from AI Feedback (RLAIF) is an approach replaces human feedback with AI-generated feedback to train a reward model and optimize an RL (Reinforcement learning) policy for alignment Bai et al. (2022).
@@
Challenges Despite their potential, scalable oversight methods face important challenges. A primary concern is the exploitability of helper systems, which are integral to these methodologies. For instance, helper systems used in training and alignment can themselves be exploited by the primary AI they are meant to align, as observed in adversarial vulnerabilities and reward hacking AdamGleave (2023); Everitt et al. (2023). Additionally, adversarial robustness remains a persistent issue, as even state-of-the-art systems are susceptible to adversarial attacks, raising concerns about their behavior under worst-case scenarios AdamGleave (2023).

### v3

@@ Abstract
The emergence of large language models (LLMs) has sparkedthe discussion on Artificial Superintelligence (ASI), a hypothetical AI system surpassing human intelligence. Though ASI is still hypothetical and far from current AI capabilities, existing alignment methods struggle to guide such advanced AI ensure its safety in the future. It is essential to discuss the alignment of such AI now. Superalignment, the alignment of AI at superhuman levels of capability systems with human values and safety requirements, aims to address two primary goals: scalability in supervision to provide high-quality guidance signals and robust governance to ensure alignment with human values. In this survey, we review the original scalable oversight problem and corresponding methods and potential solutions for superalignment. Specifically, we introduce the challenges and limitations of current alignment paradigms in addressing the superalignment problem. Then we review scalable oversight methods for superalignment. Finally, we discuss the key challenges and propose pathways for the safe and continual improvement of future AI systems. By comprehensively reviewing the current literature, our goal is provide a systematical introduction of existing methods, analyze their strengths and limitations, and discuss potential future directions.
@@ 1 Introduction
Figure 1: Challenges from the perspectives of supervision and governance. While supervision perspective focuses on providing high-quality guidance signals for enhancing system competence, governance perspective emphasizes aligning the behavior of advanced aI with human values to prevent harmful outcomes.
@@
Language models have shown improved capabilities as their scale increases, a phenomenon known as emergent abilities Wei et al. (2022a); Brown et al. (2020); Kaplan et al. (2020); Nam et al. (2024); Hu et al. (2024). This scaling has led to the development of large language models (LLMs), including proprietary models such as GPT-o1 OpenAI (2024), Claude 3.5 Anthropic (2024), and Gemini Ultra Team et al. (2024), as well as open-source models like Llama 3.2 Meta (2024), Mixtral Mistral (2024); Jiang et al. (2024), Qwen 2.5 Team (2024) and various LLM families.
@@
The emergent abilities of LLMs have enabled significant advancements across a wide range of tasks. These include natural language understanding Brown et al. (2020), reasoning Wei et al. (2022b); Rein et al. (2024), code generation Chen et al. (2021); Liu et al. (2024), and multilingual translation Shi et al. (2022). Furthermore, LLMs have demonstrated surprising breakthroughs, such as partially passing the Turing test Jones and Bergen (2024) and achieving high accuracy on hard mathematical problems Rein et al. (2024). Recent research indicates that these advancements in LLMs have sparked discussions about the transition from artificial narrow intelligence (ANI) to the possibility of artificial general intelligence (AGI), though the latter is still far from being realized. This also makes the idea of artificial superintelligence (ASI)  Pohl (2015); Batin et al. (2017) a theoretical concept worth discussing. ANI refers to AI systems currently in use, which are designed to perform specific tasks, such as LLMs for healthcare Huang et al. (2024) and software engineering Fan et al. (2023), but they lack the general cognitive abilities of humans. In contrast, AGI represents an AI model theoretical that shows human-level capabilities, including reasoning, learning, and adaptability across various domains Bubeck et al. (2023); Fei et al. (2022); Goertzel (2014); Pohl (2015); Batin et al. (2017). The development in LLM, particularly their ability to generalize knowledge and exhibit emergent behaviors, has led to the possibility of achieving AGI in the future Bubeck et al. (2023). This development also draws attention to ASI, a hypothetical future stage of AI where machines would not only hold human-level intelligence in all areas but also have advanced cognitive functions and sophisticated thinking abilities Mucci and Stryker (2023); Hughes et al. (2024). Considering the potential to achieve ASI in the future, it is essential to envision, prepare for ASI and explore how to align it with human values to avoid potential negative impact on the society.
@@
However, the development of ASI poses challenges, particularly due to the lack of guidance signals when humans cannot supervise AI systems with labeled data. Figure 1 illustrates the challenges in scaling high-quality guidance signals. Superalignment is defined as “the process of supervising, controlling, and governing artificial superintelligence systems” Jonker and McGrath (2024); Jan Leike (2023) and encompasses two sub-goals from the perspectives of supervision and governance. Achieving these goals requires an iterative and co-evolving process.
@@
From a supervision perspective, the goal is to build high-quality guidance signals to enhance the model’s helpfulness. Traditional training methods, such as reinforcement learning with human feedback (RLHF) Ouyang et al. (2022), face scalability issues once model begin to surpass human intelligence. This creates a critical bottleneck, the inability to provide high-quality guidance signals that improve the system’s helpfulness or competence. From a governance perspective, the goal is to effectively regulate the model’s behavior to ensure it remains harmless and aligned with human values. Without robust governance, ASI could behave in unpredictable, harmful, or even catastrophic. For instance, poorly aligned ASI systems might pursue goals that inadvertently disempower humanity or create existential risks. To address this, it is essential not only to constrain AI behaviors but also to ensure that these systems operate in a manner that is inherently safe and aligned with ethical principles. This Superalignment process involves developing methods to guide advanced AI systems, ensuring they remain beneficial and safe, maximizing benefits for humanity.
@@
Current alignment paradigms, such as reinforcement learning with human feedback (RLHF) Christiano et al. (2017); Ouyang et al. (2022); Gulcehre et al. (2023), supervised fine-tuning (SFT) Zhou et al. (2023); Rafailov et al. (2023), and in-context learning (ICL) Gou et al. (2024); Xu et al. (2023), have shown promise in aligning LLMs with human values and goals Wang et al. (2024). RLHF leverages human feedback to train a reward model, which generates rewards that help align the language model with human preferences. In contrast, SFT eliminates the reinforcement reward model and directly fine-tunes the language model using human feedback. Meanwhile, ICL aligns the output of LLMs with human values by constraining their responses during inference without the need for additional training. Scalable oversight, the original concept which led to the superalignment, aims to develop scalable, high-quality supervision signals that can guide AI systems beyond human capabilities while ensuring alignment with human values and goals Amodei et al. (2016); Taylor et al. (2016); Olah (2016). However, despite advancements in RLHF, SFT, and ICL, these paradigms struggle to achieve the goal of superalignment.
@@
From a supervision perspective, RLHF is limited by the scalability and reliability of human feedback. As models grow in complexity, collecting feedback that is accurate, consistent, and comprehensive becomes increasingly difficult, leaving gaps in guidance. Similarly, SFT faces challenges in curating and annotating datasets that are sufficiently diverse and representative to fine-tune ASI effectively. ICL, while bypassing training entirely, relies on carefully crafted prompts during inference, which are inherently constrained in scalability and expressiveness. From a governance perspective, RLHF and SFT rely heavily on the quality and intent of human supervision, raising concerns about biases and inconsistencies in feedback. The potential for misaligned human oversight poses risks, particularly as models exceed human capabilities. ICL, on the other hand, requires governance structures to ensure that prompt constraints align with broader human values, a process that is difficult to standardize and enforce. Additionally, the risk of deception by advanced AI systems further complicates governance Yang et al. (2024a). Thus, the lack of scalable, high-quality, and diverse supervision signals remains a bottleneck in achieving superalignment. To tackle superalignment challenges, tailored methods must be designed. As models keep advancing, it is crucial to enhance their capabilities during development while steering their values and goals to maximize benefits for human society.
@@
Our objective is to provide an overview of the superalignment problem, including its origins, historical context, and definition, as well as the current state of superalignment research and potential solutions. In Section 2, we introduce the concept of different AI types, the corresponding challenges, and the limitations of current alignment paradigms in addressing the superalignment problem. In Section 3, we analyze current scalable oversight techniques and discuss the specific challenges associated with each method. In Section 4, we summarize the key challenges of scalable oversight and explore potential pathways to address these challenges.
@@ 2.1 Terminology
Artificial Narrow Intelligence (ANI)
ANI or “weak AI,” refers to AI systems designed to excel at a single specific task $T_{0}$. These systems operate equal or below human-level performance in the task and lack generalization capabilities. Formally, This can be formalized as $\mathcal{A}_{\text{ANI}}(T_{0})\leq\mathcal{H}(T_{0})$, where $\mathcal{A}_{\text{ANI}}(T_{0})$ represents the performance of the ANI system on task $T_{0}$, and $\mathcal{H}(T_{0})$represents human-level performance for the same task and domain (e.g., translation).
However, for certain tasks, such as chess and Go, modern ANI systems have surpassed human-level performance $\mathcal{A}_{\text{ANI}}(T_{0})\textgreater\mathcal{H}(T_{0})$ Mucci and Stryker (2023); Silver et al. (2016). Despite their impressive performance in these specific areas, ANI systems are limited in their ability to perform other tasks ($T_{1:\infty}$), where $\mathcal{A}_{\text{ANI}}(T_{i})<\mathcal{H}(T_{i}),\quad\forall T_{i}\neq T_{0}$.
Artificial General Intelligence (AGI)
AGI refers to a theoretical leap beyond ANI. AGI systems would possess human-level intelligence and show general-purpose capabilities across a wide range of tasks and domains Bubeck et al. (2023); Fei et al. (2022); Goertzel (2014); Pohl (2015); Batin et al. (2017). AGI systems generalize knowledge and adapt to new challenges. This can be formalized as $\mathcal{A}_{\text{AGI}}(T_{i})=\mathcal{H}(T_{i}),\quad\forall\,T_{i}$, where $\mathcal{A}_{\text{AGI}}(T_{i})$ represents the AGI system’s performance on task $T_{i}$ with a potentially infinite set of tasks $i\to\infty$, and $\mathcal{H}(T_{i})$ represents human-level performance in the same tasks and domains. Although current LLMs mark significant progress beyond ANI, they remain far from reaching true AGI.
Artificial Superintelligence (ASI)
ASI refers to AI systems that surpass human intelligence ($\mathcal{H}$) in all tasks and domains with exceptional thinking skills Mucci and Stryker (2023); Hughes et al. (2024). ASI systems demonstrate superior reasoning, creativity, and adaptability, formally described as $\mathcal{A}_{\text{ASI}}(T_{i})\gg\mathcal{H}(T_{i}),\quad\forall\,T_{i}$, where $\gg$ indicates that the ASI system outperforms human intelligence in quality of results.
ASI is currently far from realization, but discussing how to address the challenges of superalignment is both timely and critical for mitigating catastrophic risks Pueyo (2018). Proactive preparation is essential, as the rapid development of AI could outpace our ability to respond, leaving no room for correction. Furthermore, supervision and governance must co-evolve as interactive processes, enhancing AI capabilities while simultaneously mitigating associated risks.
@@ 2.2 Superalignment
Scalable oversight is formally defined as “the process of ensuring that a given AI system adheres to aspects of its objectives that are too costly or impractical to evaluate frequently during training” Amodei et al. (2016). This concept shows the need for methods that provide reliable, high-quality supervision signals, even providing such signals is highly expensive or even infeasible as the tasks become too complicated. Superalignment, as a specific case of scalable oversight, focuses on aligning AI with human values while also enhancing its capability, where such off-the-shelf supervision signals are unavailable.
Formally, superalignment is defined as “the process of supervising, controlling, and governing artificial superintelligence systems” Jonker and McGrath (2024) with the goal of developing methods that provide high-quality supervision signals that guide its development and behavior, even in the absence of direct human oversight or traditional feedback mechanisms such as human-labeled data or RLHF Amodei et al. (2016); Christiano et al. (2018). As AI capabilities grow, these traditional methods become increasingly insufficient due to the escalating cost and expertise required for human supervision.
Since the emergence of large-scale models and the increase in their capabilities, superalignment has gained significant attention as a means to ensure that AI systems remain aligned with human values while maintaining its capability. Scalable oversight techniques, such as sandwiching and weak-to-strong generalization (W2SG), have emerged as promising approaches for superalignment.
@@ 2.3.2 Methods
Scalable oversight has emerged as a promising solution to superalignment. It enables the efficient alignment of AI systems with complex objectives that would otherwise be too costly or difficult to evaluate frequently. Various scalable oversight techniques have been proposed, including iterated distillation and amplification (IDA) Christiano et al. (2018); Charbel-Raphaël (2023), recursive reward modeling (RRM) Leike et al. (2018), and cooperative inverse reinforcement learning Hadfield-Menell et al. (2016). These approaches serve as the basis for current scalable oversight methods, such as weak-to-strong generalization, debate, reinforcement learning from AI feedback, and sandwiching Burns et al. (2023); Irving et al. (2018); Bai et al. (2022); Bowman et al. (2022). Figure 3 illustrates the current scalable oversight paradigms and corresponding concrete methodologies.
@@
  • Reinforcement Learning from AI Feedback (RLAIF) is an approach replaces human feedback with AI-generated feedback to train a reward model and optimize an RL policy model for alignment Bai et al. (2022).
@@
Challenges Despite their potential, scalable oversight methods face important challenges. A primary concern is the exploitability of helper systems, which are integral to these methodologies. For instance, helper systems used in training can themselves be exploited by the primary AI they are meant to align, as observed in adversarial vulnerabilities and reward hacking AdamGleave (2023); Everitt et al. (2023). Additionally, adversarial robustness remains a persistent issue, as even state-of-the-art systems are susceptible to adversarial attacks, raising concerns about their behavior under worst-case scenarios AdamGleave (2023).